OpenStudy (anonymous):

The length of a rectangle is 2 more than twice the width. Find the dimensions of the rectangle if the perimeter is 76. If x represents the width of the rectangle, then which expression represents the perimeter? 2x + 2 4x + 4 6x + 4

OpenStudy (kainui):

Try to make it into an equation, make the best guess you can and I'll help you fix it up until you get there. Just take it one word at a time. Call things you don't know as variables, like length can be L or Y or whatever. Eventually you should end up with some things that look like an equation.

OpenStudy (solomonzelman):

I was answering your question, why did you leave?

OpenStudy (solomonzelman):

Continuing from what we left of, l-2=2w lw=76 l-2=2w ----> l=2w + 2 lw=76 substitute, w(2w + 2)=76 2w^2+2w-76=0 divide the entire equation by 2 w^2+w-32=0 can you solve this?